Consider the following time series

Period Time series

1 48

2 41

3 37

4 32

5 36

6 31

7 43

The four-period centered moving average to time period 3 is

1. 37

2. 36:67

3. 35

4. 38

5. 35:25


1
Expert's answer
2021-08-26T13:04:21-0400

The 4-period moving average is found by counting the first four entries and finding the average.

"First \\
=\frac{48+41+37+32}{4}\\
=\frac{79}{2}\\
=39.5"


"Second\\
= \frac{41+37+32+36}{4}\\
= \frac{73}{2}\\
=36.5"

The centered

"= \frac{39.5+36.5}{2}\\
=38"

Fourth option is correct